Assume that you just won $1 million in the lottery. Ignoring taxes, how much of the...

  1. Assume that you just won $1 million in the lottery. Ignoring taxes, how much of the $1million would you spend and how much would you save? Calculate your own MPC and MPS based on your answers.
  2. Assume everyone in this class gets a $1,000 at the end of the semester.  Further assume each student consumes $750 of their bonus. Calculate the multiplier.  If there are 30 students in the class, by how much will R-GDP increase?
Answer #1

Out of $1 million lottery, I would spend $0.6 million and save $0.4 million

This means MPC = 0.6 and MPS = 0.4

----

Student MPC = Change in C / Change in Y = 750/1000 = 0.75

Multiplier = 1/(1-MPC) = 1/(1-0.75) = 4

Increase in RGDP of one student = 4 x $1000 = $12,000

Increase in RGDP of 30 students = 30 x $12000 = $360,000